Use log4^3=0.7925 and log4^4=1 to approximate the value of the expression log4^384
Show all work please and explain if you could

Answer by Alan3354(69443)

Use log4^3=0.7925 and log4^4=1 to approximate the value of the expression log4^384
--------------
You mean log base 4, not exponents.
^ is exponentiation.
----
384 = 3*128 = 3*4^3*4^(1/2)
log(384) = log(3) + 3*log(4) + (1/2)*log(4) with any consistent base
= 0.7925 + 3 + 0.5
= 4.2925